# Fine-Grained Safety Neurons with Training-Free Continual Projection to Reduce LLM Fine Tuning Risks

### v3

@@ Introduction
Table 3: Evaluating Safety and Utility under Multi-Dimensional Continual FGSN Projections.
Maintaining utility. FGSN achieves AlpacaEval Winrates of 54.61% and 54.52% on Llama-3-8B and Qwen-2.5-7B fine-tuned models, respectively, surpassing SafeLoRA (47.37%, 50.98%) and Wanda (54.15%, 51.40%). This indicates that FGSN preserves semantic understanding and response quality without significant degradation. Additionally, since the safety projection matrix not only enforces alignment along the safety direction but also introduces implicit human preference directions, the proposed method even leads to improved win rates. On the GSM8K mathematical reasoning task, FGSN achieves an accuracy of 53.20%, outperforming SafeLoRA (52.20%) and Wanda (52.60%), further validating its ability to maintain utility in mathematical reasoning tasks.
Minimal parameter editing. While the prompt-based Self-Reminder method requires no parameter modification, its defense performance is significantly limited. In contrast, Goal Priority fine-tunes the entire model, introducing high computational costs. Compared to other parameter-editing approaches such as SafeLoRA and Wanda, FGSN makes substantially fewer changes—only (Llama-3-8B: 5.38% and Qwen-2.5-7B:4.67%), resulting in minimal disruption to model utility. Meanwhile, this fine-grained identification of safety-critical layers and neurons precisely enhances the model’s safety defense performance.
Performance on Continual Safety Concerns
Safety Dimension Generalization.
To address the increasingly complex safety challenges faced by LLMs, we evaluate the proposed Continual FGSN method using multiple risk categories from the BeaverTails safety dataset (Ji et al. 2023). As shown in Tab.3, the horizontal axis represents models obtained after continual projections along safety directions targeting different dimensions, while the vertical axis reports harmfulness scores across various safety data. Results demonstrate that after the first projection using only Universal Safety data, the harmfulness scores across all evaluated dimensions decrease significantly(e.g., GPT-4o score drops from 2.04 to 1.18 on Child Abuse, and from 2.36 to 1.56 on Privacy). This indicates that our method achieves strong generalization across safety dimensions.
@@
Figure 3: Visualization of safe neuron selection across different safety dimensions.
Continual reduction in harmfulness.
@@
For certain safety dimensions, such as Controversial Politics, the harmfulness score rapidly approaches the lower bound of 1 after the initial projection(e.g., GPT-4o: 1.10 → 1.02), indicating near-optimal safety performance. In contrast, other dimensions still exhibit room for improvement. Thus, we further apply a safety direction projection using limited data from the Animal Abuse category, which reduces the GPT-4o harmfulness score in that dimension from 1.64 to 1.32. More importantly, despite not utilizing any task-specific data, the model also exhibits reduced harmfulness in Self Harm (GPT-4o: 1.58 → 1.20), Terrorism (GPT-4o: 1.74 → 1.18), and Privacy (GPT-4o: 1.56 → 1.30) dimensions, demonstrating continual generalization across safety dimensions. Subsequently, we perform continual FGSN projections using data from Child Abuse and Terrorism. The model achieves further safety improvements on these respective dimensions, with the overall harmfulness score judged by GPT decreasing to an average value of 1.15. Furthermore, it maintains the safety gains on earlier dimensions, confirming that our method supports continual learning without forgetting.
Continued Preservation of Utility.
@@
Visualization of Safety Neurons across Dimensions.
@@
Fig. 3 visualizes the distribution of safety neurons selected from Universal Safety data and those identified from various safety dimensions. Orange-red nodes represent safety neurons that overlap. The results show that, for each dimension introduced during continual learning, the majority of selected safety neurons overlap with those from the universal safety projection. Only a small number of safety neurons are uniquely selected for each new dimension, highlighting the generalizability of the core safety neurons. Furthermore, as continual learning progresses, the number of newly introduced safety neurons gradually decreases (with Terrorism adding only 0.73% of parameter modifications). This indicates that our method continues to improve safety while gradually reducing parameter modifications, leading to a stable model state. Therefore, FGSN does not introduce significant interference with the model’s generalizability.
Figure 4: Effect of different safety layer selections on model safety. In both LLaMA3.1-8B-Instruct and Qwen2.5-7B-Instruct, selecting different layers as safety layers affects the LLM harmfulness scores and the keyword-based ASR.
Ablation Study
@@
To investigate the advantages of the proposed fine-grained safety neuron selection combined with safety layers, we divide the models (LLaMA3.1-8B-Instruct and Qwen2.5-7B-Instruct) into several layer intervals. Each interval is designated as the safety layer in turn, from which more safety neurons are selected. We then evaluate the performance of the modified models on two metrics: LLM harmfulness score and keywords ASR, as shown in Fig. 4.
@@
Results on LLaMA3.1-8B-Instruct model shown in Fig. 4a indicate that replacing the safety layer in shallow layers (layers 1–5 or 5–10) does not significantly reduce the harmfulness scores or the ASR. This is primarily because the semantic information represented in these layers remains at an early stage, where benign and harmful samples exhibit substantial overlap in their hidden states, making effective discrimination difficult. Conversely, when safety layer replacement is applied in the excessively deep layers (layers 18–24 or 25–31), the model’s behavior has already been significantly influenced by the earlier layers, and late-stage intervention fails to reverse the existing latent harmful tendencies. Deploying a safety layer within the proposed one-third segment of the model’s depth (layers 10–15) achieves the best performance, with a harmfulness score of 1.02 and an ASR of 14%. This placement effectively balances semantic representation and behavioral control, enabling more precise localization of safety neurons and thereby enhancing both safety and efficacy. Consistent results were also obtained in the Qwen2.5-7B-Instruct model.
Conclusion
@@
In this work, we introduce Fine-Grained Safety Neurons with Training-Free Continual Projection (FGSN), a novel defense framework designed to mitigate safety risks introduced by fine-tuning LLMs. By leveraging safe-relevant neuron localization based safety-critical layers, FGSN achieves precise identification of safety neurons while minimizing interference with general tasks. Through sparse, training-free projections along safety directions, our approach not only enhances defense effectiveness under static conditions but also supports continual adaptation to emerging safety threats. Extensive experiments across multiple models and benchmarks validate that FGSN consistently improves safety with minimal parameter edits and negligible degradation in utility. Furthermore, we demonstrate the generalizability of safety neurons across risk dimensions and the sustainability of incremental safety enhancements. However, its ability to continual learning across more complex and diverse safety dimensions remains an area for further investigation. Overall, our work highlights the importance of fine-grained model introspection for robust safety alignment and opens new directions for scalable, maintenance-free safety defenses in the evolving LLM landscape.

### v4

@@ 4.1 Experimental Settings
Table 3: Evaluating Safety and Utility under Multi-Dimensional Continual MSCP Projections.
Evaluation Metrics. We follow Chao et al. (2024) to evaluate the proposed method from safety and utility. For safety, we adopt both LLM-based safety judger and keyword-based attack success rate (ASR). The LLM judger score ranges from 1 to 10, with higher scores indicating greater unsafe behavior. For utility, we use the AlpacaEval Li et al. (2023) semantic evaluation benchmark to compute the WinRate between the original LoRA-finetuned model and the MSCP-modified model on the Alpaca dataset. We use accuracy to evaluate performance on the GSM8K dataset.
4.2 Performance on Static Safety Concerns
Enhancing safety defenses. In terms of safety judged by closed-source GPT-4o and open-source Llama3.1-405B, MSCP consistently achieves the lowest harmfulness scores across both Alpaca-finetuned models (1.02 / 1.27 on Llama-3-8B and 1.37 / 1.36 on Qwen-2.5-7B), clearly outperforming alternative methods such as Goal Priority ($\geq$2.10), Self Reminder ($\geq$2.01), and SafeLoRA ($\geq$1.51) as Tab. 1. Moreover, MSCP achieves the lowest keyword ASR on Alpaca-finetuned models (14% on both Llama and Qwen), demonstrating superior robustness compared to Goal Priority (35%, 30%) and Wanda (30%, 16%). These results indicate MSCP’s effectiveness in mitigating a broad spectrum of safety risks. Additionally, on the GSM8K-finetuned model, MSCP maintains strong performance, with GPT-4o and Llama3.1-405B harmfulness scores of 1.94 and 1.93, respectively—outperforming all other defense-capable baselines as Tab. 2.
@@ 4.2 Performance on Static Safety Concerns
Figure 3: a):Visualization of safe neuron selection across different safety dimensions.b) and c): In both LLaMA3.1-8B-Instruct and Qwen2.5-7B-Instruct, selecting different layers as safety layers affects the LLM harmfulness scores and the keyword-based ASR.
Maintaining utility. MSCP achieves AlpacaEval Winrates of 54.61% and 54.52% on Llama-3-8B and Qwen-2.5-7B fine-tuned models, respectively, surpassing SafeLoRA (47.37%, 50.98%) and Wanda (54.15%, 51.40%). This indicates that MSCP preserves semantic understanding and response quality without significant degradation. Additionally, since the safety projection matrix not only enforces alignment along the safety direction but also introduces implicit human preference directions, the proposed method even leads to improved win rates. On the GSM8K mathematical reasoning task, MSCP achieves an accuracy of 53.20%, outperforming SafeLoRA (52.20%) and Wanda (52.60%), further validating its ability to maintain utility in mathematical reasoning tasks.
Minimal parameter editing. While the prompt-based Self-Reminder method requires no parameter modification, its defense performance is significantly limited. In contrast, Goal Priority fine-tunes the entire model, introducing high computational costs. Compared to other parameter-editing approaches such as SafeLoRA and Wanda, MSCP makes substantially fewer changes—only (Llama-3-8B: 5.38% and Qwen-2.5-7B:4.67%), resulting in minimal disruption to model utility. Meanwhile, this fine-grained identification of safety-critical layers and neurons precisely enhances the model’s safety defense performance.
4.3 Performance on Continual Safety Concerns
4.3.1 Safety Dimension Generalization.
To address the increasingly complex safety challenges faced by LLMs, we evaluate the proposed Continual MSCP method using multiple risk categories from the BeaverTails safety dataset Ji et al. (2023). As shown in Tab.3, the horizontal axis represents models obtained after continual projections along safety directions targeting different dimensions, while the vertical axis reports harmfulness scores across various safety data. Results demonstrate that after the first projection using only Universal Safety data, the harmfulness scores across all evaluated dimensions decrease significantly(e.g., GPT-4o score drops from 2.04 to 1.18 on Child Abuse, and from 2.36 to 1.56 on Privacy). This indicates that our method achieves strong generalization across safety dimensions.
4.3.2 Continual reduction in harmfulness.
@@ 4.3 Performance on Continual Safety Concerns
For certain safety dimensions, such as Controversial Politics, the harmfulness score rapidly approaches the lower bound of 1 after the initial projection(e.g., GPT-4o: 1.10 → 1.02), indicating near-optimal safety performance. In contrast, other dimensions still exhibit room for improvement. Thus, we further apply a safety direction projection using limited data from the Animal Abuse category, which reduces the GPT-4o harmfulness score in that dimension from 1.64 to 1.32. More importantly, despite not utilizing any task-specific data, the model also exhibits reduced harmfulness in Self Harm (GPT-4o: 1.58 → 1.20), Terrorism (GPT-4o: 1.74 → 1.18), and Privacy (GPT-4o: 1.56 → 1.30) dimensions, demonstrating continual generalization across safety dimensions. Subsequently, we perform continual MSCP projections using data from Child Abuse and Terrorism. The model achieves further safety improvements on these dimensions, with the overall harmfulness score judged by GPT decreasing to an average value of 1.15. Furthermore, it maintains the safety gains on earlier dimensions, confirming that our method supports continual learning without forgetting.
4.3.3 Continued Preservation of Utility.
@@
4.3.4 Visualization of Safety Neurons across Dimensions.
@@
Fig. 3a visualizes the distribution of safety neurons selected from Universal Safety data and those identified from various safety dimensions. Orange-red nodes represent safety neurons that overlap. The results show that, for each dimension introduced during continual learning, the majority of selected safety neurons overlap with those from the universal safety projection. Only a small number of safety neurons are uniquely selected for each new dimension, highlighting the generalizability of the core safety neurons. Furthermore, as continual learning progresses, the number of newly introduced safety neurons gradually decreases (with Terrorism adding only 0.73% of parameter modifications). This indicates that our method continues to improve safety while gradually reducing parameter modifications, leading to a stable model state. Therefore, MSCP does not introduce significant interference with the model’s generalizability.
4.4 Ablation Study
@@
To investigate the advantages of the proposed multi-Level safety localization, we divide the models into several layer intervals. Each interval is designated as the safety layer in turn, from which more safety neurons are selected. We then evaluate the performance of the modified models on two metrics: LLM harmfulness score and keywords ASR, as Fig. 3.
@@
Results on LLaMA3.1-8B-Instruct model shown in Fig. 3b indicate that replacing the safety layer in shallow layers (layers 1–5 or 5–10) does not significantly reduce the harmfulness scores or the ASR. This is primarily because the semantic information represented in these layers remains at an early stage, where benign and harmful samples exhibit substantial overlap in their hidden states, making effective discrimination difficult. Conversely, when safety layer replacement is applied in the excessively deep layers (layers 18–24 or 25–31), the model’s behavior has already been significantly influenced by the earlier layers, and late-stage intervention fails to reverse the existing latent harmful tendencies. Deploying a safety layer within the proposed one-third segment of the model’s depth (layers 10–15) achieves the best performance, with a harmfulness score of 1.02 and an ASR of 14%. This placement effectively balances semantic representation and behavioral control, enabling more precise localization of safety neurons and thereby enhancing both safety and efficacy. Consistent results were also obtained in the Qwen2.5-7B-Instruct model.
5 Conclusion
@@
In this work, we introduce Fine-Grained Safety Neurons with Training-Free Continual Projection (MSCP), a novel defense framework designed to mitigate safety risks introduced by fine-tuning LLMs. By characterizing the coordinated interactions between global and localized safety representations across multiple levels, we precisely identify highly sparse neuron clusters that govern safety-sensitive behaviors, while minimizing interference with general-purpose task performance. Through sparse, training-free projections along safety directions, our approach not only enhances defense effectiveness under static conditions but also supports continual adaptation to emerging safety threats. Extensive experiments across multiple models and benchmarks validate that MSCP consistently improves safety with minimal parameter edits and negligible degradation in utility. However, its ability to continual learning across more complex and diverse safety dimensions remains an area for further investigation. Overall, our work highlights the importance of fine-grained model introspection for robust safety alignment and opens new directions for scalable, maintenance-free safety defenses in the evolving LLM landscape.